# Step1X-Edit: A Practical Framework for General Image Editing

### v1

@@ Abstract
In recent years, image editing technology has witnessed remarkable and rapid development. The recent unveiling of cutting-edge multimodal models such as GPT-4o and Gemini2 Flash has introduced highly promising image editing capabilities. These models demonstrate an impressive aptitude for fulfilling a vast majority of user - driven editing requirements, marking a significant advancement in the field of image manipulation. However, there are still a large gap between the open-sourced algorithm with these closed-source models. Thus, in this paper, we target to release a state-of-the-art image editing model, called Step1X-Edit, which can provide comparable performance against the closed-source models like GPT-4o and Gemini2 Flash. More specifically, we adopt the Multimodal LLM to process the reference image and user’s editing instruction. A latent embedding has been extracted and integrated with a diffusion image decoder to obtain the target image. To train the model, we build a data generation pipeline to produce a high-quality dataset. For evaluation, we develop the GEdit-Bench, a novel benchmark rooted in real-world user instructions. Experimental results on GEdit-Bench demonstrate that Step1X-Edit outperforms existing open-source baselines by a substantial margin and approaches the performance of leading proprietary models, thereby making significant contributions to the field of image editing.
@@ 1 Introduction
Image editing with natural language instructions has become an increasingly important task in vision-language research. It offers intuitive interaction for end users while posing unique technical challenges: understanding nuanced semantics, precisely localizing regions to edit, and preserving image fidelity. While diffusion models have dramatically improved image generation quality, the existing design by integrating text encoder, e.g., CLIP [33] and T5 [34], with diffusion transformer often struggles with following editing instruction to maintain alignment between input image and edit instruction, especially when edit instructions are subtle or compositional.
@@ 2 Related Work
2.1 Controllable Image Generation and Editing
Autoregressive (AR) models have been actively studied for controllable image generation and editing by modeling images as sequences of discrete tokens. Works such as ControlAR [23], ControlVAR [21], and CAR [54] incorporate spatial and pixel-level guidance—like edges, segmentation masks, or depth maps—into the decoding process, enabling more structured and localized control. Training-Free VAR [48] enables editing without inversion through token-level distribution caching and adaptive masking, while M2M [39], MSGNet [9], and MVG [37] extend AR-based generation to multi-object, temporal, and multimodal scenarios. However, due to reliance on discrete visual tokens and sequence length limits, AR models often struggle to produce high-resolution and photorealistic results, especially in complex scenes.
Diffusion models, by contrast, have become the dominant approach for high-fidelity image synthesis, with strong capabilities in photorealism, structural consistency, and diversity. Beginning with DDPM [16] and DDIM [42], and further advanced by Latent Diffusion [38, 32, 62], these models operate in latent spaces for improved scalability. ControlNet [58] and T2I-Adapter [28] inject spatial or task-specific control into the generation process, while InstructPix2Pix [8], Pix2Pix-Zero [31], and UniDiffuser [4] support instruction-based or zero-shot editing without retraining. Despite these strengths, diffusion models often depend on static prompts or predefined conditions and lack the capacity for multi-turn reasoning or flexible language alignment, limiting their use in open-ended editing scenarios.
These limitations have led to growing interest in unified image editing frameworks that combine the symbolic control of AR models with the generative fidelity of diffusion. Such models aim to tightly couple instruction understanding, spatial reasoning, and photorealistic synthesis within a single architecture, offering more flexible, general, and user-controllable editing capabilities. We discuss this direction further in the next section on Unified Image Editing Models.
@@ 2.2 Unified Instruction-based Image Editing Models
Unified image editing models aim to bridge semantic instruction understanding with precise visual manipulation in a single, coherent framework. Early approaches often rely on modular designs, where MLLMs generate textual prompts or spatial instructions to guide diffusion models, as seen in Prompt-to-Prompt [15], InstructEdit [45], and BrushEdit [22]. InstructPix2Pix [8] trains a conditional diffusion model using synthetic instruction-image pairs, while MagicBrush [57] improves real-world usability through high-quality human annotations. To tackle data scarcity and alignment issues, SeedEdit [41] distills text-to-image models into consistent editing models via causal diffusion and iterative self-alignment. Recent works such as SmartEdit [17], X2I [26], and RPG [53] enhance interaction between language and vision by integrating multimodal attention and reasoning strategies. AnyEdit [55] introduces task-aware routing and visual prompt projection within a unified diffusion model, while OmniGen [51] adopts a single transformer backbone to jointly encode text and images, removing reliance on external encoders. More generally, models such as Gemini [14] and GPT-4o [29] demonstrate strong visual fluency through joint vision-language training, showing promising capabilities in understanding and generating consistent, context-aware images. Collectively, these developments reflect a shift from loosely coupled systems toward tightly integrated, instruction-driven editing frameworks.
@@ 3.1.2 Caption Strategy
Cost-Efficient Pipeline: To control annotation costs while maintaining quality, we first use GPT-4o to perform the aforementioned annotation procedures. The annotated results are then used to fine-tune our in-house Step1o model, which is employed to scale up annotation for larger datasets in a more cost-effective manner.
@@
Figure 3: Data Construction Pipeline and Sub-Task Distribution.
@@ 3.2 Our Method
As illustrated in Fig. 4, our algorithm involves three key components: a Multimedia Large Language Model (MLLM), a connector module, and a Diffusion in Transformer (DiT). The input editing instruction, accompanied by the reference image, is first introduced to the MLLM; e.g., Qwen-VL [3] (hereafter abbreviated as Qwen). In conjunction with a system prefix, these inputs are jointly processed through a single forward pass of the MLLM, enabling the model to capture the semantic relationships between the instruction and the visual content. To isolate and emphasize the semantic elements relevant to the editing task, we selectively discard the token embeddings associated with the prefix. This filtering process retains only the token embeddings that directly align with the edit instruction, ensuring that subsequent processing focuses precisely on the user-specified editing requirements.
@@
The extracted embeddings are then fed into a lightweight connector module, such as the token refiner [25, 19]. This module restructures the embeddings into a more compact textual feature representation. The refined feature subsequently substitutes the text embedding that was initially generated by the T5 [35] encoder in the downstream DiT network, e.g., FLUX [6]. Furthermore, we calculate the mean of all output embeddings from Qwen. This mean value is then projected through a linear layer, resulting in the generation of a global visual guidance vector. By doing so, the image editing network can leverage Qwen’s enhanced semantic comprehension capabilities, enabling more accurate and context-aware editing operations.
To effectively train the Token Refiner and enable rich cross-modal conditioning, we draw inspiration from the token concatenation mechanism introduced in FLUX-Fill [7]. The key idea is to enhance the model’s ability to reason over contrastive visual contexts. During training, both a target image and a reference image are fed into the system. The target image is first encoded by a VAE encoder, followed by the addition of Gaussian noise to promote generalization. The resulting latent is then linearly projected into an image token representation. In contrast, the reference image is encoded without noise and projected similarly. These two sets of image tokens are concatenated along the token length dimension, forming a fused feature of doubled token length, which is used as the final visual input.
The model is trained in a joint learning setup, where the connector and the downstream DiT are optimized simultaneously. We initialize both components with pretrained weights from our in-house Qwen and DiT text-to-image model, enabling better convergence and performance transfer. The learning rate is set to $1e^{-5}$, balancing training stability and convergence speed.
By combining structured language guidance, token-level visual conditioning, and strong pretrained backbones within a unified framework, our method significantly boosts the system’s capability to perform high-fidelity, semantically aligned image edits across a diverse range of user instructions.

### v2

@@ Abstract
In recent years, image editing technology has witnessed remarkable and rapid development. The recent unveiling of cutting-edge multimodal models such as GPT-4o and Gemini2 Flash has presented highly promising image editing capabilities. These models demonstrate an impressive aptitude for fulfilling a vast majority of user-driven editing requirements, marking a significant advancement in the field of image manipulation. However, there is still a large gap between the open-source algorithm with these closed-source models. To this end, we introduce a state-of-the-art image editing model, Step1X-Edit, which aims to provide comparable performance against the closed-source models like GPT-4o and Gemini2 Flash. More specifically, we adopt the Multimodal LLM to process the reference image and the user’s editing instruction. A latent embedding has been extracted and integrated with a diffusion image decoder to obtain the target image. To train this model, we build a data generation pipeline covering 11 editing tasks to produce a high-quality dataset. For evaluation, we develop the GEdit-Bench, a novel benchmark rooted in real-world user instructions. Experimental results on GEdit-Bench demonstrate that Step1X-Edit outperforms existing open-source baselines by a substantial margin and approaches the performance of leading proprietary models, thereby making significant contributions to the field of image editing.
@@ 1 Introduction
Image editing with natural language instructions has become an increasingly important task in vision-language research. It offers intuitive interaction for end users while posing unique technical challenges: understanding nuanced semantics, precisely localizing regions to edit, and preserving image fidelity. While diffusion models [37, 5, 10, 36] have dramatically improved image generation quality, the existing design by integrating text encoder, e.g., CLIP [39] and T5 [40], with diffusion transformer often struggles with following editing instruction to maintain alignment between input image and edit instruction, especially when edit instructions are subtle or compositional.
@@ 2 Related Work
2.1 Controllable Image Generation and Edit
Autoregressive (AR) models have been actively studied for controllable image generation and editing by modeling images as sequences of discrete tokens. Works such as ControlAR [26], ControlVAR [24], and CAR [59] incorporate spatial and pixel-level guidance—such as edges, segmentation masks, and depth maps—into the decoding process, enabling localized and structured control. Extensions like Training-Free VAR [53], M2M [43], and Instruct-imagen [18] further improve editing flexibility and broaden application scenarios. UniFluid [11] explores unified autoregressive generation and understanding with continuous visual tokens. However, due to reliance on discrete tokens and sequence length constraints, AR models often struggle to produce high-resolution and photorealistic results, especially in complex scenes.
Diffusion models have become the dominant approach for high-fidelity image synthesis, offering strong capabilities in photorealism, structural consistency, and diversity. Beginning with DDPM [17] and DDIM [46], and further advanced by Latent Diffusion [42, 37], diffusion models operate in latent spaces for improved scalability. With the introduction of DiT architectures [36], diffusion models have made significant strides in generalization, image quality, and knowledge capacity, becoming the predominant architecture in modern image generation [1, 5]. Based on the above text2image models, ControlNet [63], T2I-Adapter [32], and OmniControl [57] inject spatial or task-specific control into the generation process. BrushNet [21], PowerPaint [68], and FLUX-Fill [6] further improve inpainting quality and versatility. Despite these advances, diffusion models often rely on static prompts or fixed conditions and lack the capacity for multi-turn reasoning or flexible language alignment, limiting their application in open-ended editing scenarios.
These limitations have led to growing interest in unified image editing frameworks that combine the symbolic control of AR models with the generative fidelity of diffusion. Such models aim to tightly couple instruction understanding, spatial reasoning, and photorealistic synthesis within a single architecture, offering more flexible, general, and user-controllable editing capabilities.
@@ 2.2 Unified Instruction-based Image Editing Models
Unified instruction-based image editing models aim to bridge semantic instruction understanding and precise visual manipulation. Early approaches such as Prompt-to-Prompt [16], InstructEdit [50], InstructPix2Pix [7], MagicBrush [62], and BrushEdit [25] adopt modular pipelines where MLLMs generate prompts, spatial cues, or synthetic instruction-image pairs to guide diffusion-based editing.
Recent works move toward tighter integration between instruction and generation. SmartEdit [19], X2I [30], RPG [58], AnyEdit [60], and UltraEdit [66] enhance multimodal interaction and instruction fidelity through improved model architectures, task-aware routing, and fine-grained editing capabilities. Meanwhile, unified generation and editing frameworks such as OmniGen [56], ACE [15], ACE++[31], and Lumina-OmniLV[38] consolidate diverse visual tasks under a single architecture. Methods like OminiControl [48], Qwen2VL-Flux [28], and MetaQueries [34] further explore efficient control integration and direct latent-level fusion between MLLMs and diffusion decoders.
More generally, models such as Gemini [14] and GPT-4o [33] demonstrate strong visual fluency through joint vision-language training, showing promising capabilities in understanding and generating consistent, context-aware images. Collectively, these developments reflect a shift from loosely coupled systems toward tightly integrated, instruction-driven editing frameworks.
@@ 3.1.1 Data Pipeline
Figure 3: Data Construction Pipeline and Sub-Task Distribution.
@@ 3.2 Our Method
As illustrated in Fig. 4, our algorithm involves three key components: a Multimedia Large Language Model (MLLM), a connector module, and a Diffusion in Transformer (DiT) [36]. The input editing instruction, accompanied by the reference image, is first introduced to the MLLM; e.g., Qwen-VL [3] (hereafter abbreviated as Qwen). In conjunction with a system prefix, these inputs are jointly processed through a single forward pass of the MLLM, enabling the model to capture the semantic relationships between the instruction and the visual content. To isolate and emphasize the semantic elements relevant to the editing task, we selectively discard the token embeddings associated with the prefix. This filtering process retains only the token embeddings that directly align with the edit information, ensuring that subsequent processing focuses precisely on the editing requirements.
@@
The extracted embeddings are then fed into a lightweight connector module, such as the token refiner [29, 22]. This module restructures the embeddings into a more compact textual feature representation, which is subsequently used as the text embedding input to the downstream DiT network. Furthermore, we calculate the mean of valid embeddings from Qwen. This mean value is then projected through a linear layer, generating a global visual guidance vector. By doing so, the image editing network can leverage Qwen’s enhanced semantic comprehension capabilities, enabling more accurate and context-aware editing operations.
To effectively train the Token Refiner and enable rich cross-modal conditioning, we carefully design our feature aggregation strategy. Compared to approaches such as FLUX-Fill [6], which uses channel concatenation, and methods like SeedEdit [45], which introduce additional causal self-attention mechanisms, we follow OminiControl [48] and adopt token concatenation to better balance responsiveness to editing instructions with the preservation of fine-grained image details. During training, the reference image is encoded by a VAE encoder, and its latent features are linearly projected into reference image tokens. After processing, the image tokens are concatenated with noise image tokens along the token length dimension to construct the final visual input.
Recent research has also explored alternative approaches to enhance cross-model understanding by leveraging MLLMs. For instance, Qwen2VL-FLUX [28] introduces a pioneering approach by replacing the traditional T5 [40] text encoder in DiT-based text-to-image models with MLLMs to enhance multi-modal understanding and generation. However, it still retains T5 [40] for text encoding, which limits its ability to perform comprehensive cross-modal reasoning. In contrast, approaches like DreamEngine [8] leverages Qwen to align image and text modalities, with their features serving as external conditional inputs for SD3.5 [1] in the image synthesis. This method establishes a shared representation space that facilitates a more coherent process of generation and understanding. Nevertheless, in DreamEngine, the challenge remains in fully capturing the fine-grained details of reference images using only MLLM features.
Compared to these approaches, our model not only retains cross-modal understanding but also enhances the extraction of image details. By combining structured visual-language guidance, detailed visual condition, and strong pretrained backbones within a unified framework, our method significantly boosts the system’s capability to perform high-fidelity, semantically aligned image edits across a diverse range of user instructions. During training, we jointly optimize the connector and the downstream DiT using only the diffusion loss, following the rectified flow [10] formulation. Moreover, our method ensures stable training without relying on the mask loss trick, which distinguishes it from OmniGen [56]. The learning rate is fixed at $1e^{-5}$ to ensure a good trade-off between training stability and convergence speed.